# Classification of thin Jordan schemes

### v1

@@ 3.1. Non-regular thin Jordan schemes
From (4), it follows that without loss of generality we may assume that $\mathcal{G}_{1}=\mathcal{G}_{2}$, and that $\varphi$ is the identity-map. From now on we make this assumption.
A immediate consequence from (3) is, that
$\forall_{A\in\mathcal{G}_{1}},\,\forall_{C\in\mathcal{H}}\,:\,AC=CA,$ (5)
since $A=\varphi(A)=C^{-1}AC$.
Note now that
$\forall_{A\in\mathcal{G}_{1}},\,\forall_{C\in\mathcal{H}}\,:\,AC\in\mathcal{H}.$ (6)
To see this we simply compute for given $A\in\mathcal{G}_{1}$ and $C\in\mathcal{H}$:
$\begin{bmatrix}A&O\\[2.84526pt] O&A\end{bmatrix}\star\begin{bmatrix}O&C\\[2.84526pt] O&O\end{bmatrix}=\frac{1}{2}\begin{bmatrix}O&AC+CA\\[2.84526pt] O&O\end{bmatrix}=\begin{bmatrix}O&AC\\[2.84526pt] O&O\end{bmatrix}$.
Next, we show that $\mathcal{G}_{1}$ is an abelian group. To this end, let $C\in\mathcal{H}$, and let $A,A^{\prime}\in\mathcal{G}_{1}$. Then
$AA^{\prime}C=A(A^{\prime}C)=(A^{\prime}C)A=A^{\prime}AC.$
Thus, $AA^{\prime}=A^{\prime}A$, whence $A\star A^{\prime}=AA^{\prime}$, and $\mathcal{G}_{1}$ is indeed an abelian group.
It remains to show that $\mathcal{H}=\mathcal{K}=\mathcal{G}_{1}$. For this, let $C=(c_{i,j})$ be the unique element of $\mathcal{H}$ with $c_{1,1}=1$ (in other words, $C\vec{e}_{1}=\vec{e}_{1}$; this exists because $\sum\mathcal{H}=J$). Let $k\in\{1,\dots,n\}$ and let $A_{k}=(a^{(k)}_{i,j})$ be the unique element of $\mathcal{G}_{1}$ with $a^{(k)}_{1,k}=1$ (in other words, $A_{k}\vec{e}_{1}=\vec{e_{k}}$; this exists because $\sum\mathcal{G}_{1}=J$).
Since $A_{k}C\colon\vec{e}_{1}\mapsto\vec{e}_{k}$, and $CA_{k}\colon\vec{e}_{1}\mapsto C\vec{e}_{k}$, and since $A_{k}C=CA_{k}$ (by Claim 7), it follows that $C=I$. But now it follows from Claim 8 that $\mathcal{H}=\{AC\mid A\in\mathcal{G}_{1}\}=\mathcal{G}_{1}$.
As $\mathcal{K}=\mathcal{H}^{-1}$, we have also $\mathcal{K}=\mathcal{G}_{1}$. By now it is clear that $\mathcal{A}=\mathcal{J}(\mathcal{G}_{1})$. ∎
@@ 3.2. Regular thin Jordan schemes
In this subsection, we assume that $(\Omega,S)$ is a regular thin JS, that is, $S$ is a set of permutations on $\Omega$. The set $S$ contains the identity permutation $1_{\Omega}$ and is closed when transposed, that is, $s^{t}\in S$ for every $s\in S$. Notice that $s^{t}$ is a permutation inverse to $s$, that is, $s^{-1}=s^{t}$.
Since $S$ is a partition of $\Omega^{2}$ into a disjoint union of permutations, for every pair $(\alpha,\beta)\in\Omega^{2}$ the basic relation $S(\alpha,\beta)\in S$ is the unique permutation that maps $\alpha$ to $\beta$. From the definition of a JS it follows that $S(\alpha,\alpha)=1_{\Omega}$ and $S(\alpha,\beta)=S(\beta,\alpha)^{-1}$. The same definition implies that the adjacency matrix $(S(\alpha,\beta))_{\alpha,\beta\in\Omega}$ is a Latin square. Pick an arbitrary point $\omega_{0}\in\Omega$ and define a binary operation $\diamond$ on $S$ as follows $a\diamond b=S(\omega_{0},a(b(\omega_{0})))$, that is, a triple $a,b,c\in S$ satisfies $a\diamond b=c$ iff $a(b(\omega_{0}))=c(\omega_{0})$.
@@
It follows from the definition that the product $\diamond$ depends a choice of ”base point” $\omega_{0}$. It could be that a choice of a different base point will lead to a non-isomorphic magma.
@@
  1. (1)
    $1_{\Omega}$ is neutral element of $(S,\diamond)$;
  2. (2)
@@
(1) We have $a\diamond 1_{\Omega}=S(\omega_{0},a(1_{\Omega}(\omega_{0})))=S(\omega_{0},a(\omega_{0}))=a$. Similarly, $1_{\Omega}\diamond a=a$.
(2) This part follows from the identity $a^{t}(a(\omega_{0}))=\omega_{0}=1_{\Omega}(\omega_{0})$.
(3) $a\diamond b=c\Leftrightarrow a(b(\omega_{0}))=c(\omega_{0})\Leftrightarrow b(\omega_{0})=a^{t}(c(\omega_{0}))\Leftrightarrow a^{t}\diamond c=b$.
(4) We have $x(a(\omega_{0}))=b(\omega_{0})$. Therefore, $x=S(a(\omega_{0}),b(\omega_{0}))$. ∎
It follows from the above statement that $(S,\diamond)$ is a loop with left inverse property. It is well-known that each element $a\in S$ of a loop determines two permutations $\ell_{a}$ and $r_{a}$ on $S$, defined naturally by $\ell_{a}(x)=a\diamond x$, and $r_{a}(x)=x\diamond a$. From (3) it follows that $\ell_{a^{t}}\ell_{a}=1_{S}$, that is, $\ell_{a^{t}}=\ell_{a}^{-1}$. From (3), it follows that the left division in the loop $(S,\diamond)$ is $a^{t}\diamond b$. The right division will be denoted as $b/a$, that is, $(b/a)\diamond a=b$. Later we will show that $b/a=b\diamond a^{t}$.
@@
  1. (1)
    $\sigma$ is a bijection;
  2. (2)
    $\sigma^{-1}a\sigma=\ell_{a}$, for each $a\in S$;
  3. (3)
@@
(1) If $s\neq t\in S$, then $s\cap t=\emptyset$, implying that $\sigma(s)=s(\omega_{0})\neq t(\omega_{0})=\sigma(t)$. Thus, $\sigma$ is injective. To prove surjectivity, pick an arbitrary $\omega\in\Omega$ and set $s:=S(\omega_{0},\omega)$. Then $\sigma(s)=s(\omega_{0})=\omega$.
(2) We have to show that $a\sigma=\sigma\ell_{a}$ holds for each $a\in S$. The latter equality is equivalent to the following one $\forall_{s\in S}:\ a(\sigma(s))=\sigma(\ell_{a}(s))$. Now $a(\sigma(s))=a(s(\omega_{0}))=(a\diamond s)(\omega_{0})=\sigma(\ell_{a}(s))$.
(3) We have to show that $S(\sigma(u),\sigma(v))=1_{S}(v/u)=v/u$ holds for all $u,v\in S$. Denote $w:=v/u$. Then $w\diamond u=v$ implying $w(u(\omega_{0}))=v(\omega_{0})$. Hence $w=S(\sigma(u),\sigma(v))$. ∎
The Proposition above implies that the permutations $\{\ell_{s}\}_{s\in S}$ form a Jordan scheme on $S$ which is isomorphic to the original one: $(\Omega,S)$. Therefore $\forall_{u,v\in S}\ \underline{\ell_{u}}\cdot\underline{\ell_{v}}+\underline{\ell_{v}}\cdot\underline{\ell_{u}}=\underline{\ell_{u\diamond v}}+\underline{\ell_{v\diamond u}}$, or equivalently,
@@
This equality may be rephrased as
@@
As an immediate consequence of the above property we conclude that $(S,\diamond)$ has the left alternative property: $u\diamond(u\diamond v)=(u\diamond u)\diamond v$.
For a pair $u,v\in S$ we set $A_{u,v}:=\{w\in S\mid u\diamond(v\diamond w)=(u\diamond v)\diamond w\}$. Notice that $1_{\Omega}\in A_{u,v}$. It follows from (8) that $A_{u,v}=A_{v,u}$.
@@ 3.2.1. Properties of $\diamond$
We will show here that $(S,\diamond)$ is a loop. We start with the following statement.
@@
    If $u\diamond v=v\diamond u$, then $\ell_{u}\ell_{v}=\ell_{v}\ell_{u}=\ell_{u\diamond v}$ and $A_{u,v}=S$;
@@
Part (a) follows from (8).
Part (b). It follows from $u\diamond v\neq v\diamond u$ that $\ell_{u\diamond v}\cap\ell_{v\diamond u}=\emptyset$. Therefore $(u\diamond v)\diamond w\neq(u\diamond v)\diamond w$. Now (8) implies the claim. ∎
@@
Let $u,v\in S$ and let $A:=A_{u,v}$, $\bar{A}:=S\setminus A$. Then the following are true:
@@
    $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$ (in particular, $\ell_{u\diamond v}\subseteq\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}$);
@@
If $u$ and $v$ commute, then we are done by Proposition 3.7(a). So, for the rest of the proof we assume that $u,v$ do not commute.
Part (a). By (7) $\ell_{u}\ell_{v}\cup\ell_{v}\ell_{v}=\ell_{u\diamond v}\cup\ell_{v\diamond u}$.
By (8) $(u\diamond v)\diamond w\in\{u\diamond(v\diamond w),v\diamond(u\diamond w)\}$. More precisely,
@@
Equivalently, $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$.
Part (b). It follows from $u^{t}\diamond(u\diamond v)=v$ and from Part (a) that $\ell_{v}\subseteq\ell_{u^{t}}\ell_{u\diamond v}\cup\ell_{u\diamond v}\ell_{u^{t}}$. Taking into account that $\ell_{u^{t}}=\ell_{u}^{-1}$ we obtain
$\ell_{v}\subseteq\ell_{u}^{-1}(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\cup(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\ell_{u}^{-1}.$
After simplifications we obtain
@@
Let us count the intersection of $\ell_{v}$ with each summand of (9):
@@
This implies that $(\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}\ell_{u}^{-1})\cap\ell_{v}=\ell_{v}$ which, in turn, yields $\ell_{u}1_{\bar{A}}\ell_{u}^{-1}=1_{\bar{A}}$, or equivalently, $\ell_{u}(\bar{A})=\bar{A}$. The equality $\ell_{v}(\bar{A})=\bar{A}$ follows from $u\leftrightarrow v$ symmetry. ∎
It follows from Proposition 3.8(b) that $u\diamond A_{u,v}=A_{u,v}$, and that $v\diamond A_{u,v}=A_{u,v}$. Together with $1_{\Omega}\in A_{u,v}$ we obtain that the subloop of $(S,\diamond)$ generated by $u$ and $v$ is contained in $A_{u,v}$. In particular, $u,v\in A_{u,v}$.
@@
It follows from $u\diamond A_{u,v}=A_{u,v}$ that $u^{t}\diamond(u\diamond A_{u,v})=u^{t}\diamond A_{u,v}$. By Proposition 3.5 $u^{t}\diamond(u\diamond A_{u,v})=A_{u,v}$, Hence $u^{t}\diamond A_{u,v}=A_{u,v}$. Analogously, $v^{t}\diamond A_{u,v}=A_{u,v}$. Now it follows from $1_{\Omega}\in A_{u,v}$ that $u^{t}\diamond v^{t}\in A_{u,v}=A_{v,u}$. Therefore $(v\diamond u)\diamond(u^{t}\diamond v^{t})=v\diamond(u\diamond(u^{t}\diamond v^{t}))=v\diamond v^{t}=1_{\Omega}$. Together with $(v\diamond u)\diamond(v\diamond u)^{t}=1_{\Omega}$ we obtain $(v\diamond u)^{t}=(u^{t}\diamond v^{t})$. ∎
@@
The loop $(S,\diamond)$ is alternative, that is for all $u,v\in S$ we have
@@
Part (L) follows from equation (8). Parts (R) and (M) follow from the inclusion $u\in A_{u,v}=A_{v,u}$. ∎
@@
A loop is Moufang iff it is simultaneously left and right Bol loop. We start by showing that $(S,\diamond)$ is a left Bol loop. To this end we need to prove the validity of the identity
@@
This identity is equivalent to $\ell_{y}\ell_{z}\ell_{y}=\ell_{y\diamond(z\diamond y)}$. To see that the latter identity holds, we note that there exists some $w\in S$ such that $\ell_{y}\ell_{z}\ell_{y}=\ell_{w}$ (see Proposition 5.1 in the Appendix). Further we argue that since $(S,\{\ell_{s}\}_{s\in S})$ is a Jordan scheme (in particular we use that in a Jordan scheme two relations are either equal or disjoint), we have that $\ell_{w}=\ell_{y\diamond(z\diamond y)}$ if and only if $\ell_{w}(1_{S})=\ell_{y\diamond(z\diamond y)}(1_{S})$. It remains to compute:
@@
Hence, identity (10) holds and, $(S,\diamond)$ is indeed a left Bol loop.
To see that $(S,\diamond)$ is also a right Bol loop we need to prove the validity of the identity
@@
Again we note that this is equivalent to the identity $r_{y}r_{z}r_{y}=r_{(y\diamond z)\diamond y}$. From Proposition 3.9 it follows that $r_{z}=\tau\ell_{z^{t}}\tau$ where $\tau\in\mathsf{Sym}(S)$ is the transposing involution on $S$, i.e. $\tau(x)=x^{t}$. Since $(S,\diamond)$ is a left Bol loop, we have $\ell_{y^{t}}\ell_{z^{t}}\ell_{y^{t}}=\ell_{y^{t}\diamond(z^{t}\diamond y^{t})}$ (see above). If we conjugate both sides of this identity by $\tau$, then we obtain on one hand
@@
and on the other hand
$\tau\ell_{y^{t}\diamond(z^{t}\diamond y^{t})}\tau=\tau\ell_{((y\diamond z)\diamond y)^{t}}\tau=r_{(y\diamond z)\diamond y},$
as desired. Thus Identity (11) holds, too, and $(S,\diamond)$ is a right Bol loop.∎
@@
If the Jordan scheme $(\Omega,S)$ is thin and symmetric, then $S$ is an abelian $2$-group and $(\Omega,S)$ is a symmetric association scheme.
@@
Since $a^{t}=a$, each element of the loop $(S,\diamond)$ has order two. It’s well-known that a Moufang loop of exponent two is a group. Hence $(S,\diamond)$ is an elementary abelian $2$-group. ∎

### v2

@@ 3.1. Nonregular thin Jordan schemes
By (4) we may now assume that $\mathcal{G}_{1}=\mathcal{G}_{2}$ and moreover that $\varphi$ is the identity-map. Henceforth, we make these assumptions.
Since $A=\varphi(A)=C^{-1}AC$, an immediate consequence of (3) is
$\forall_{A\in\mathcal{G}_{1}},\,\forall_{C\in\mathcal{H}}\,:\,AC=CA.$ (5)
For $A\in\mathcal{G}_{1}$ and $C\in\mathcal{H}$, we next compute
$\begin{bmatrix}A&O\\[2.84526pt] O&A\end{bmatrix}\star\begin{bmatrix}O&C\\[2.84526pt] O&O\end{bmatrix}=\frac{1}{2}\begin{bmatrix}O&AC+CA\\[2.84526pt] O&O\end{bmatrix}=\begin{bmatrix}O&AC\\[2.84526pt] O&O\end{bmatrix}$
which proves
$\forall_{A\in\mathcal{G}_{1}},\,\forall_{C\in\mathcal{H}}\,:\,AC\in\mathcal{H}.$
We next claim $\mathcal{G}_{1}$ is an abelian group. To this end, let $C\in\mathcal{H}$ and let $A,A^{\prime}\in\mathcal{G}_{1}$. Then
$AA^{\prime}C=A(A^{\prime}C)=(A^{\prime}C)A=A^{\prime}AC\implies AA^{\prime}=A^{\prime}A\implies A\star A^{\prime}=AA^{\prime},$
and the claim is proved.
Finally, it remains to show that $\mathcal{H}=\mathcal{K}=\mathcal{G}_{1}$. For this purpose, let $C=(c_{i,j})$ be the unique element of $\mathcal{H}$ with $c_{1,1}=1$, i.e. $C\vec{e}_{1}=\vec{e}_{1}$. (Note that this element exists because $\sum\mathcal{H}=J$.) Let $k\in\{1,\dots,n\}$ and let $A_{k}=\big(a^{(k)}_{i,j}\big)$ be the unique element of $\mathcal{G}_{1}$ with $a^{(k)}_{1,k}=1$, i.e. $A_{k}\vec{e}_{1}=\vec{e_{k}}$. (This element exists due to $\sum\mathcal{G}_{1}=J$.)
Since $A_{k}C\colon\vec{e}_{1}\mapsto\vec{e}_{k}$, $CA_{k}\colon\vec{e}_{1}\mapsto C\vec{e}_{k}$, and $A_{k}C=CA_{k}$ by (5), it follows that $C=I$. We thereby conclude that $\mathcal{H}=\{AC\mid A\in\mathcal{G}_{1}\}=\mathcal{G}_{1}$. As $\mathcal{K}=\mathcal{H}^{-1}$, we also have $\mathcal{K}=\mathcal{G}_{1}$. But now it is clear that $\mathcal{A}=\mathcal{J}(\mathcal{G}_{1})$. ∎
@@ 3.2. Regular thin Jordan schemes
In this subsection, we assume that $(\Omega,S)$ is a regular thin Jordan scheme, that is, $S$ is a set of permutations on $\Omega$. The set $S$ contains the identity permutation $1_{\Omega}$ and is closed under transpose, i.e. $s^{t}\in S$ for every $s\in S$. In fact $s^{t}=s^{-1}$.
Since $S$ is a partition of $\Omega^{2}$ into a disjoint union of permutations, for every pair $(\alpha,\beta)\in\Omega^{2}$ the basic relation $S(\alpha,\beta)\in S$ is the unique permutation that maps $\alpha$ to $\beta$. From the definition of a Jordan scheme it follows that $S(\alpha,\alpha)=1_{\Omega}$ and $S(\alpha,\beta)=S(\beta,\alpha)^{-1}$. The same definition implies that the adjacency matrix $(S(\alpha,\beta))_{\alpha,\beta\in\Omega}$ is a Latin square. Pick an arbitrary point $\omega_{0}\in\Omega$ and define a binary operation $\diamond$ on $S$ by $a\diamond b=S(\omega_{0},a(b(\omega_{0})))$. That is, a triple $a,b,c\in S$ satisfies $a\diamond b=c$ iff $a(b(\omega_{0}))=c(\omega_{0})$.
@@
It follows from the definition that the product $\diamond$ depends on the choice $\omega_{0}$ of “base point”. It could be that a different choice of base point will yield a nonisomorphic magma.
@@
  1. (a)
    $1_{\Omega}$ is the neutral element of $(S,\diamond)$;
  2. (b)
@@
Proof of a. We have $a\diamond 1_{\Omega}=S(\omega_{0},a(1_{\Omega}(\omega_{0})))=S(\omega_{0},a(\omega_{0}))=a$. Similarly, $1_{\Omega}\diamond a=a$.
Proof of b. This follows at once from the identity $a^{t}(a(\omega_{0}))=\omega_{0}=1_{\Omega}(\omega_{0})$.
Proof of c. $a\diamond b=c\Leftrightarrow a(b(\omega_{0}))=c(\omega_{0})\Leftrightarrow b(\omega_{0})=a^{t}(c(\omega_{0}))\Leftrightarrow a^{t}\diamond c=b$.
Proof of d. We have $x(a(\omega_{0}))=b(\omega_{0})$. Therefore, $x=S(a(\omega_{0}),b(\omega_{0}))$. ∎
It follows from the above statement that $(S,\diamond)$ is a loop with the left inverse property. It is well-known that each element $a\in S$ of a loop determines two permutations $\ell_{a}$ and $r_{a}$ on $S$, defined naturally by $\ell_{a}(x)=a\diamond x$, and $r_{a}(x)=x\diamond a$. From c it follows that $\ell_{a^{t}}\ell_{a}=1_{S}$, that is, $\ell_{a^{t}}=\ell_{a}^{-1}$. It also follows from c that left division in the loop $(S,\diamond)$ is $a^{t}\diamond b$. Right division will be denoted as $b/a$, that is, $(b/a)\diamond a=b$. Later we will prove $b/a=b\diamond a^{t}$.
@@
  1. (a)
    $\sigma$ is a bijection.
  2. (b)
    $\sigma^{-1}a\sigma=\ell_{a}$ for each $a\in S$.
  3. (c)
@@
Proof of a. If $s\neq t\in S$, then $s\cap t=\emptyset$. Thus $\sigma(s)=s(\omega_{0})\neq t(\omega_{0})=\sigma(t)$ which proves $\sigma$ is injective. To prove surjectivity, pick an arbitrary $\omega\in\Omega$ and set $s:=S(\omega_{0},\omega)$. Then $\sigma(s)=s(\omega_{0})=\omega$.
Proof of b. We must show $a\sigma=\sigma\ell_{a}$ for each $a\in S$. But this is equivalent to showing $\forall_{s\in S}:\ a(\sigma(s))=\sigma(\ell_{a}(s))$. Now $a(\sigma(s))=a(s(\omega_{0}))=(a\diamond s)(\omega_{0})=\sigma(\ell_{a}(s))$.
Proof of c. We must show $S(\sigma(u),\sigma(v))=1_{S}(v/u)=v/u$ for all $u,v\in S$. Denote $w:=v/u$. Then $w\diamond u=v$ which implies $w(u(\omega_{0}))=v(\omega_{0})$. Hence $w=S(\sigma(u),\sigma(v))$. ∎
Proposition 3.6 implies that the permutations $\{\ell_{s}\}_{s\in S}$ form a Jordan scheme on $S$ isomorphic to the original $(\Omega,S)$. Therefore $\forall_{u,v\in S}\ \underline{\ell_{u}}\cdot\underline{\ell_{v}}+\underline{\ell_{v}}\cdot\underline{\ell_{u}}=\underline{\ell_{u\diamond v}}+\underline{\ell_{v\diamond u}}$, or equivalently,
@@
This latter equality may be re-expressed as
@@
As an immediate consequence we conclude that $(S,\diamond)$ has the left alternative property: $u\diamond(u\diamond v)=(u\diamond u)\diamond v$.
For each pair $u,v\in S$, we set $A_{u,v}:=\{w\in S\mid u\diamond(v\diamond w)=(u\diamond v)\diamond w\}$. Observe that $1_{\Omega}\in A_{u,v}$. It follows from (7) that $A_{u,v}=A_{v,u}$.
@@ 3.2.1. Properties of $\diamond$
We will now show that $(S,\diamond)$ is a loop. We start with the following assertion.
@@
    If $u\diamond v=v\diamond u$, then $\ell_{u}\ell_{v}=\ell_{v}\ell_{u}=\ell_{u\diamond v}$ and $A_{u,v}=S$.
@@
Proof of (a). This is an immediate consequence of (7).
Proof of (b). As $u\diamond v\neq v\diamond u$, we have $\ell_{u\diamond v}\cap\ell_{v\diamond u}=\emptyset$. But then $(u\diamond v)\diamond w\neq(v\diamond u)\diamond w$, whence the result follows once more from (7). ∎
@@
Let $u,v\in S$, and let $A:=A_{u,v}$, $\bar{A}:=S\setminus A$. Then the following hold:
@@
    $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$; in particular, $\ell_{u\diamond v}\subseteq\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}$.
@@
If $u$ and $v$ commute, then we are done by Proposition 3.7(a). So for the balance of the proof we assume $uv\neq vu$.
Proof of (a). By (6), $\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}=\ell_{u\diamond v}\cup\ell_{v\diamond u}$ while by (7), $(u\diamond v)\diamond w\in\{u\diamond(v\diamond w),v\diamond(u\diamond w)\}$. More precisely,
@@
But this is equivalent to $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$.
Proof of (b). From (a) and the fact that $u^{t}\diamond(u\diamond v)=v$, it follows that $\ell_{v}\subseteq\ell_{u^{t}}\ell_{u\diamond v}\cup\ell_{u\diamond v}\ell_{u^{t}}$. Taking into account that $\ell_{u^{t}}=\ell_{u}^{-1}$ we now obtain
$\ell_{v}\subseteq\ell_{u}^{-1}(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\cup(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\ell_{u}^{-1}$
which simplifies to
@@
Below we determine the cardinality of the intersection of $\ell_{v}$ with each set appearing in (8).
@@
Together, this implies $(\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}\ell_{u}^{-1})\cap\ell_{v}=\ell_{v}$ which in turn yields $\ell_{u}1_{\bar{A}}\ell_{u}^{-1}=1_{\bar{A}}$, or equivalently, $\ell_{u}(\bar{A})=\bar{A}$. Similarly, we obtain the equality $\ell_{v}(\bar{A})=\bar{A}$ from $u\leftrightarrow v$ symmetry. ∎
It follows from Proposition 3.8(b) that $u\diamond A_{u,v}=A_{u,v}$ and $v\diamond A_{u,v}=A_{u,v}$. Together with $1_{\Omega}\in A_{u,v}$, we see that the subloop of $(S,\diamond)$ generated by $u$ and $v$ is contained in $A_{u,v}$. In particular, $u,v\in A_{u,v}$.
@@
Since $u\diamond A_{u,v}=A_{u,v}$, it follows that $u^{t}\diamond(u\diamond A_{u,v})=u^{t}\diamond A_{u,v}$. By Proposition 3.5c, $u^{t}\diamond(u\diamond A_{u,v})=A_{u,v}$ whence $u^{t}\diamond A_{u,v}=A_{u,v}$. Analogously, $v^{t}\diamond A_{u,v}=A_{u,v}$. Now since $1_{\Omega}\in A_{u,v}$, we obtain $u^{t}\diamond v^{t}\in A_{u,v}=A_{v,u}$. Therefore $(v\diamond u)\diamond(u^{t}\diamond v^{t})=v\diamond(u\diamond(u^{t}\diamond v^{t}))=v\diamond v^{t}=1_{\Omega}$. Together with $(v\diamond u)\diamond(v\diamond u)^{t}=1_{\Omega}$, we conclude that $(v\diamond u)^{t}=u^{t}\diamond v^{t}$ as desired. ∎
@@
The loop $(S,\diamond)$ is alternative, that is to say, for all $u,v\in S$ we have
@@
Equality (L) follows from equation (7), while (R) and (M) follow from the inclusion $u\in A_{u,v}=A_{v,u}$. ∎
@@
A loop is Moufang iff it is simultaneously a left and right Bol loop. We start by showing that $(S,\diamond)$ is a left Bol loop. To this end, we must prove the identity
@@
But this identity is equivalent to $\ell_{y}\ell_{z}\ell_{y}=\ell_{y\diamond(z\diamond y)}$. To establish the latter identity, we first note that $\mathcal{J}:=\langle\underline{\ell}_{s}\rangle_{s\in S}$ is $\star$-closed. Therefore $\underline{\ell}_{y}\underline{\ell}_{z}\underline{\ell}_{y}=2(\underline{\ell}_{y}\star(\underline{\ell}_{z}\star\underline{\ell}_{y})-(\underline{\ell}_{y}\star\underline{\ell}_{y})\star\underline{\ell}_{z})\in\mathcal{J}$. This implies the permutation matrix $\underline{\ell}_{y}\underline{\ell}_{z}\underline{\ell}_{y}$ is a linear combination of pairwise disjoint permutation matrices $\underline{\ell}_{s},s\in S$. But this can only occur if $\underline{\ell}_{y}\underline{\ell}_{z}\underline{\ell}_{y}=\underline{\ell}_{w}$ for some $w\in S$, or equivalently, $\ell_{y}\ell_{z}\ell_{y}=\ell_{w}$. We further argue that since $(S,\{\ell_{s}\}_{s\in S})$ is a Jordan scheme, we have $\ell_{w}=\ell_{y\diamond(z\diamond y)}$ if and only if $\ell_{w}(1_{S})=\ell_{y\diamond(z\diamond y)}(1_{S})$. (We have here used the fact that in a Jordan scheme two relations are either equal or disjoint.) We now compute
@@
Thus identity (9) holds, and $(S,\diamond)$ is indeed a left Bol loop.
To see that $(S,\diamond)$ is a right Bol loop, we must establish the identity
@@
Similarly to the above, this identity is equivalent to $r_{y}r_{z}r_{y}=r_{(y\diamond z)\diamond y}$. From Proposition 3.9 it follows that $r_{z}=\tau\ell_{z^{t}}\tau$ where $\tau\in\mathsf{Sym}(S)$ is the involutory transpose permutation on $S$, i.e. $\tau(x)=x^{t}$. Since $(S,\diamond)$ is a left Bol loop, we have $\ell_{y^{t}}\ell_{z^{t}}\ell_{y^{t}}=\ell_{y^{t}\diamond(z^{t}\diamond y^{t})}$. Conjugating both sides of this identity by $\tau$, we obtain on one hand
@@
while on the other hand
$\tau\ell_{y^{t}\diamond(z^{t}\diamond y^{t})}\tau=\tau\ell_{((y\diamond z)\diamond y)^{t}}\tau=r_{(y\diamond z)\diamond y}.$
Thus identity (10) holds, and $(S,\diamond)$ is a right Bol loop as well.∎
@@
Let $(\Omega,S)$ be a thin symmetric Jordan scheme. Then $S$ is an abelian $2$-group and $(\Omega,S)$ is a symmetric association scheme.
@@
Since $a^{t}=a$, each element of the loop $(S,\diamond)$ has order two. It is well known that a Moufang loop of exponent two is a group. Hence $(S,\diamond)$ is an elementary abelian $2$-group. ∎